OpenStudy (anonymous):

Use implicit differentiation to find the equation of the tangent line to the curve xy^3+xy=10 at the point (5,1) . The equation of this tangent line can be written in the form y=mx+b where m is: and b is:

OpenStudy (anonymous):

first step is to find the derivative of \(y\) wrt \(x\) start with \[y^3+3x^2y'+y+xy'=0\] replace \(x\) by \(5\) and \(y\) by \(1\) and solve for \(y'\)

OpenStudy (anonymous):

y prime= -10?

OpenStudy (anonymous):

y^3 + 3y^2* y' *x + y + xy' = 0 y' = -(y^3 +y)/(x+3y^2 *x)

OpenStudy (anonymous):

-1/10

OpenStudy (anonymous):

and then whats next?

OpenStudy (anonymous):

that's your 'm' (for the tangent line) use point-slope or y=mx+b 1=(-1/10) *5 +b b=...

OpenStudy (anonymous):

3/2 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

all good? questions?

OpenStudy (anonymous):

i understand i got 3/8 but i understand what i did wrong thanks!